OpenStudy (anonymous):

can someone please take a look at these i really dont know how to explain to my kids.. i will try to download them

OpenStudy (anonymous):

take a screen shot if you can

OpenStudy (anonymous):

let x be the amount of water in ounces at the beginning

OpenStudy (anonymous):

For each step there will be a new equation

OpenStudy (anonymous):

let's use the children's initials to represent how much water is remaining after a kid gets a drink

OpenStudy (anonymous):

\[I=x-0.25x=0.75x\]

OpenStudy (anonymous):

M=I-6

OpenStudy (anonymous):

P=M-0.2M=0.8M

OpenStudy (anonymous):

E=P-P/6=P*5/6

OpenStudy (anonymous):

Y=E-0.1*E=0.9E

OpenStudy (anonymous):

S=Y=9

OpenStudy (anonymous):

Then all there is to do is a bunch of substitution starting with substituting "I" into M=I-6.

OpenStudy (anonymous):

make sure that they are all in the same current, you don't want litter and ounce mess up

OpenStudy (anonymous):

acutally not litters but the number without currency

OpenStudy (anonymous):

my kids will be confused cause this is hard for me to explain.. there are two different questions.. sorry not math smart.. :(

OpenStudy (anonymous):

it's ok, that's y we have this website

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

but that person's explaining up there might screw you up, also ur kids

OpenStudy (anonymous):

they are just in middle school.. so i need to explain it so that we all understand

OpenStudy (anonymous):

i think it should be easy, you just keep track of the numbers and change it into fraction but 6 and 9

OpenStudy (anonymous):

then add them up

OpenStudy (anonymous):

ok.. 1/4 4divided into 100 0.2 not sure and 10% not sure how to change it..

OpenStudy (anonymous):

here i got this 1/4 + 1/6 + 1/5 + 1/10 + 6 + 9 = x

OpenStudy (anonymous):

dont the bottoms all have to match

OpenStudy (anonymous):

20 ounces

OpenStudy (anonymous):

The amount of water changes during each step and is much more complicated than simply adding up the amounts at each step.

OpenStudy (anonymous):

If we work backwards we can say that 9 is 90% of what was there before Yorick drank

OpenStudy (anonymous):

\[9=0.9*Amount before Yorick\] \[\frac{9 }{ 0.9 }=Amount before Yorick\] Amount before Yorick = 10

OpenStudy (anonymous):

We can continue doing this. Now that we know the amount before Yorick we can find the amount before Elizabeth

OpenStudy (anonymous):

Amount before Yorick=5/6*Amount before Elizabeth

OpenStudy (anonymous):

10=5/6AmtBElizabeth AmtBeforeElizabeth=\[\frac{ 10 }{ 5/6 }=12\]

OpenStudy (anonymous):

so we have to work backward ?

OpenStudy (anonymous):

its easier since we don't have to substitute all the way through

OpenStudy (anonymous):

easier than what i was describing before, breaks the problem into smaller steps

OpenStudy (anonymous):

yeah, i didnt realize that it seems math is harder and harder each year

OpenStudy (anonymous):

these types of problems are simply annoying because they can involve a lot of algebra.

OpenStudy (anonymous):

my daughter got 1/2

OpenStudy (anonymous):

0.5 ounces?

OpenStudy (anonymous):

she wrote 1/2

OpenStudy (anonymous):

ok thank you for your help.. i am so confused now. :)

OpenStudy (anonymous):

=)) we all do

OpenStudy (anonymous):

1/2 just doesn't make sense unless it is in a different unit

OpenStudy (anonymous):

she is not good in math either.. she said she did it in class with other kids.. so i am not sure how this all goes

OpenStudy (anonymous):

let's start over go to the end of the list. we know that stephanie ends with 9 ounces

OpenStudy (anonymous):

and we know that the 9 ounces are 10% less than the amount before Yorick drinks

OpenStudy (anonymous):

i think you should give her the most you can do, they she will check with friends in class

OpenStudy (anonymous):

yes thank you.. this is way to confusing for me.. lol... i should have taken a math class during the summer to help

OpenStudy (anonymous):

good night